if focus is(1,2) and eq.of directrix is 3x-4y-15=0 .find equation of parabola.sir i attempted but my answer is coming wrong

Grade:11

1 Answers

vikas askiitian expert
13 years ago

eq of directrix : 3x-4y-15 = 0                    (given)

focus = (1,2)                                           (given)

perpendicular distance from directrix to any point on parabola is equals to distance bw any point on parabola to focus...

3x-4y-15/5 = {(x-1)2+(y-2)2}1/2

squaring both sides

(3x-4y-15)2 = 25{(x-1)2+(y-2)2}2

this is the required eq of parabola